OpenStudy (anonymous):

A rectangular prism has a volume of 48 ft 3. If the height of the prism is 4 ft, and the length is 3 feet, what is the width?

Mehek (mehek14):

\(A=L*W*H\\48=4*3*W\)

Mehek (mehek14):

first thing id to multiply \(4*3\)

Mehek (mehek14):

is*

OpenStudy (anonymous):

Thanks for the help!

Mehek (mehek14):

did you get the answer?

Mehek (mehek14):

\(4*3=12\\ \dfrac{48}{12}=3\) so width = 3